# ABC: Achieving Better Control of Multimodal Embeddings using VLMs

### v1

@@ 3.1 Model Design
Table 1: Comparison of retrieval performance on MSCOCO (Lin et al., 2015) and Flickr30K (Plummer et al., 2016) datasets (Karpathy split). Best performance is bold, second best is underlined. * indicates a closed-weight model.
3.2 Data and Training
To create our pretraining dataset we employ negative mining on Conceptual Captions (Sharma et al., 2018). We derive the mined negatives for our dataset as follows: (1) We do a small pretraining run using only in-batch negatives. (2) We use the resulting model to calculate similarity scores between all images and captions in our pretraining dataset. This approach avoids a circular dependence on a third-party embedding model to train our embedding model. Therefore, it is easily extensible to modalities where an existing embedding model is not publicly available. (3) To prevent our negatives from being too similar to our positive samples, we set a similarity threshold $\epsilon\in[0,1]$. We only sample negatives that have a similarity score of at most $\epsilon$ times the similarity score of the correct candidate. We randomly choose our mined hard negatives from the $100$ candidate captions below the threshold. Our approach is similar to a negative mining technique from NV-Retreiver (de Souza P. Moreira et al., 2024), a text embedding model. This results in the text candidates shown in Figure 3. The mined text negatives are clearly relevant, but the correct caption is still the best answer.
Stage 1: Pretraining with mined negatives. In our pretraining run, each batch consists of $N$ image queries (without instructions) and $M$ text candidates. We include $M-N$ mined text negatives in each batch. Therefore, each image query has $\frac{M}{N}-1$ corresponding mined negatives. Our pretraining loss function is given by Equation 2.
$-\sum_{i}^{N}\log\frac{\exp({\bm{x}}_{i}^{\top}{\bm{y}}_{i}/\tau)}{\sum_{j=1}^{N}(\exp(\frac{{\bm{x}}_{i}^{\top}{\bm{y}}_{j}}{\tau})+\sum_{k=1}^{\frac{M}{N}-1}\exp(\frac{{\bm{x}}_{i}^{\top}{\bm{n}}_{j}^{k}}{\tau}))}$ (2)
For a given query ${\bm{x}}_{i}$, ${\bm{n}}_{i}^{k}$ is its $k_{th}$ corresponding mined hard negative and and ${\bm{y}}_{i}$ is its positive caption. $\tau$ is the temperature hyperparameter used to scale the loss. ${\bm{n}}_{j}^{k}$ is a mined negative for ${\bm{x}}_{i}$ when $i=j$, otherwise it acts as a regular in-batch negative. The embeddings for ${\bm{x}}_{i}$, ${\bm{y}}_{i}$ and ${\bm{n}}^{k}_{j}$ are unit normalized before loss is computed. We scale the loss by $\frac{1}{N}$, the number of image queries in the batch.
Instruction design. To create our instruction fine-tuning dataset we use Visual Genome (Krishna et al., 2016), a dataset comprised of images with captioned bounding boxes. When choosing which bounding boxes to use for each image, we filter by the size of the bounding box (width * height). We exclude the $5$ largest bounding boxes, as we find they often do not represent specific objects or aspects of the scene. For each image, we randomly choose $4$ bounding boxes and their respective captions. We then prompt GPT-4o (OpenAI, 2024) to create instructions corresponding to each bounding box caption. We sample multiple captions from each image so that they can be used as negatives for each other during instruction fine-tuning. This ensures that each image query has multiple realistic captions in its batch. Therefore, utilizing the instruction is required to choose the correct text candidate unambiguously. We provide the prompt and generation settings used to create our instruction fine-tuning dataset in Appendix A.
Stage 2: Instruction Fine-tuning. Following the standard VLM instruction format, we insert the instruction tokens after the image tokens in our prompt template:
[…]
We instruction fine-tune using exclusively in-batch negatives. However, we group all queries that contain the same image into the same batch. This ensures that queries always have multiple relevant text candidates. We do not back-propagate through text candidate embeddings, only the queries containing an image and instruction. All text candidates are embedded using the frozen model from stage 1. This ensures that the embeddings of our image-text queries share the same features as the pretrained model. We can easily alternate between embedding with or without instructions by disabling the instruction fine-tuned LoRA adapter.
@@ 3.2 Data and Training
ImageNet-1K (Russakovsky et al., 2015) 55.8 63.5 45.4 58.3 48.0 10.3 49.1 71.2
HatefulMemes (Kiela et al., 2021) 51.1 51.7 47.2 56.4 49.0 49.6 51.0 52.1
VOC2007 (Everingham et al., ) 50.7 52.4 64.3 66.2 51.6 52.1 74.6 81.4
SUN397 (Xiao et al., 2010) 43.4 68.8 39.6 63.2 57.0 34.5 60.1 71.8
Place365 (López-Cifuentes et al., 2020) 28.5 37.8 20.0 36.5 31.5 21.5 35.3 40.7
ImageNet-A (Djolonga et al., 2020) 25.5 14.2 42.6 9.8 8.0 3.2 31.6 49.4
ImageNet-R (Hendrycks et al., 2021) 75.6 83.0 75.0 66.2 70.9 39.7 66.2 86.8
ObjectNet (Barbu et al., 2019) 43.4 51.4 40.3 32.2 31.6 20.6 49.2 67.7
Country-211 (Radford et al., 2021) 19.2 16.8 14.2 11.3 6.2 2.5 9.3 18.5
@@
VQA (10 tasks)
OK-VQA (Marino et al., 2019) 7.5 11.5 2.4 25.4 12.7 8.7 28.0 48.1
A-OKVQA (Schwenk et al., 2022) 3.8 3.3 1.5 8.8 2.9 3.2 11.6 37.3
DocVQA (Mathew et al., 2021b) 4.0 5.3 4.2 6.2 3.0 2.6 12.6 28.5
InfographicsVQA (Mathew et al., 2021a) 4.6 4.6 2.7 4.6 5.9 2.0 10.6 7.9
ChartQA (Masry et al., 2022) 1.4 1.5 3.0 1.6 0.9 0.5 2.4 11.7
Visual7W (Zhu et al., 2016) 4.0 2.6 1.2 14.5 2.5 1.3 9.0 25.6
ScienceQA (Lu et al., 2022) 9.4 10.2 7.9 12.8 5.2 6.8 23.3 26.3
VizWiz (Gurari et al., 2018) 8.2 6.6 2.3 24.3 1.7 4.0 25.9 29.4
GQA (Hudson & Manning, 2019) 41.3 52.5 57.5 48.8 43.5 9.7 41.3 60.1
TextVQA(Singh et al., 2019) 7.0 10.9 1.0 15.1 4.6 3.3 18.9 35.4
All VQA 9.1 10.9 8.4 16.2 8.3 4.2 18.4 31.0
@@
Table 2: Zero-shot classification and VQA results on MMEB (Jiang et al., 2024b). We compare with pretrained CLIP models (Radford et al., 2021; Cherti et al., 2023a), instruction finetuned models derived from CLIP (Wei et al., 2023; Zhang et al., 2024) and other LLM backbone approaches (Li et al., 2023; Zhou et al., 2024).
4 Experiments
Benchmarks. We evaluate two aspects of the model. First, we assess the overall quality of the embeddings output by the model. We evaluate our model on a variety of zero-shot retrieval, classification and VQA tasks (Section 4.1). We further demonstrate that scaling VLM encoder resolution during test time correlates with a significant improvement on certain benchmarks (Section 4.2). Secondly, we measure our model’s ability accomplish tasks that are ambiguous without natural language instructions. To measure this, we construct CtrlBench, a benchmark where the model must use both the image and instruction to retrieve the most appropriate caption (Section 4.4).
Ablations. We ablate over several model design and training decisions. We find that our mined negative pretraining regime increases performance and stabilizes loss dynamics during pretraining (Section 4.3). We evaluate our choice of VLM backbone and its effect on representation quality (Section 4.5). Lastly, we provide ablations over adapter and attention configuration in Appendix B.
Training settings. We use Qwen2-VL-7B  (Wang et al., 2024b) as our VLM backbone for all experiments. We train our negative mining model using only in-batch negatives with a batch size of $256$ for $1000$ steps. We set $\epsilon=0.95$ and sample 7 mined negatives for each image. We pretrain using batches of $512$ image queries and $4096$ text candidates sharded across 8 NVIDIA A100-SXM4-80GB GPUs (Qu et al., 2021) for 4000 steps. We use a LoRA adapter with a rank of $64$ and an alpha of $128$. We limit the number of tokens output by the vision encoder to $512$ during training. For our optimizer, we use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $4\times 10^{-5}$, betas of $0.9$ and $0.999$ and a weight decay of $10^{-3}$. We warm-up for 3% of training steps and initialize the temperature $\tau$ as $7\times 10^{-2}$. In our instruction fine-tuning stage, we use a lower rank LoRA adapter with rank and alpha of $16$ and $32$, respectively. Our instruction fine-tuning stage can be short, as our VLM backbone is already instruction fine-tuned. Therefore, we only instruction fine-tune for $100$ steps. Each batch contains 128 unique images, with each image appearing four times, paired with a different instruction and a corresponding positive text candidate.
Modifications to save VRAM. Due to our use of LoRA (Hu et al., 2021), the VRAM requirements for gradients and optimizer state is relatively low. Consequentially, the model activations used by autograd for the backward pass account for most of VRAM used during training. To address this, we make aggressive use of activation checkpointing, recomputing the activations of each decoder block during the backward pass. Furthermore, we modify the VLM backbone to skip the logits and cross-entropy loss computation. With the large vocabulary size of modern LLMs, the output layer has become the most memory-intensive layer (Wijmans et al., 2024). As we only use the hidden state for calculating embeddings, the logits tensor is unnecessary. We find that this simple change saves up to 11 GB of memory per device, allowing us to use to a significantly larger batch size.
Figure 4: (Left) Pixel distributions of benchmarks. (Right) Scaling the number of pixels (tokens) in the vision encoder.
4.1 Zero-shot Evaluations
Table 1 demonstrates the retrieval capabilities of our model on MSCOCO (Lin et al., 2015) and Flickr30K (Plummer et al., 2016). Our model demonstrates impressive image-to-text retrieval capabilities, achieving competitive performance to models that have been contrastively trained with hundreds of GPUs and massive batch sizes (Sun et al., 2023; Cherti et al., 2023a). We achieve the best performance in MSCOCO image-to-text retrieval. Comparatively, our text-to-image retrieval is weaker. This follows from us training use image-to-text negatives, not text-to-image negatives.
@@ 4.1 Zero-shot Evaluations
To benchmark our model’s zero-shot image classification and VQA abilities, we use MMEB (Jiang et al., 2024b), a collection of multimodal embedding tasks. We note that the classification labels in MMEB are short, often only one or two words. This is problematic for image-captioning models that have been trained on full sentences (Yu et al., 2022). To alleviate this issue, we use Radford et al. (2021)’s technique of embedding classification labels in a sentence template. We use “A photo of a {label}.” as our template for all classification evaluations. We use instructions for VQA but not for image-level retrieval or classification, as we find that instructions are unnecessary for these tasks.
Our zero-shot classification and VQA results (Table 2) are very strong compared to other multimodal embedding models. We average $11.2\%$ better in classification and $12.6\%$ better in VQA than the next-best model. We find that plausible negatives are crucial during instruction fine-tuning, not just pretraining. In just 100 training steps using instructions that include negatives, our model surpasses MMRet (Zhou et al., 2024) on VQA, which was trained with millions of instructions. Interestingly, CLIP derivatives fine-tuned with instructions such as MagicLens (Zhang et al., 2024) and UniIR (Wei et al., 2023) do not perform significantly better than their respective baselines, despite MMEB providing instructions for each task. This evidences that encoder-only embedding models struggle to effectively utilize natural language instructions (Jiang et al., 2024b).
4.2 Image Resolution during Test Time
@@
Qwen2-VL-7B (Wang et al., 2024b) supports image inputs with variable resolution. This allows the user to effectively trade-off image resolution with inference speed by adjusting the number of tokens output by the vision encoder. We examine how this trade-off influences embedding quality for image classification. We find that performance on certain tasks, like ObjectNet (Barbu et al., 2019) and ImageNet-A (Djolonga et al., 2020), strongly correlate with the resolution used in the VLM vision encoder (Figure 4). On average, ObjectNet images have $13$ times more pixels than those from ImageNet-1K. When the number of tokens produced by the vision encoder is scaled up, we see a large improvement on ObjectNet, with accuracy increasing by 23.4%. However, lower resolution benchmarks like ImageNet-1K do not benefit nearly as much from scaling resolution. Notably, a smartphone camera takes photos with significantly higher resolution than images in ImageNet-1K or ObjectNet, by default (Apple Inc., ). This motivates the need for benchmarks containing larger resolution images. Existing low-resolution benchmarks may under predict the capabilities of models that are able to natively utilize high resolution images. Furthermore, we find that correctly classifying the “natural adversarial examples” of ImageNet-A (Djolonga et al., 2020) is largely a function of resolution. Our accuracy on ImageNet-A increases from only $12\%$ to $47.5\%$ simply by scaling the resolution used by the vision encoder during evaluation.
@@ 4.3 Temperature and Loss Dynamics
The setting of the temperature hyperparameter ($\tau$) is known to be crucial for contrastive training (Jia et al., 2021). We find that a poor treatment of the $\tau$ parameter is the easiest way to lose model performance. In particular, optimizing $\tau$ throughout the pretraining process is crucial. In shorter runs like training our negative mining model, we find that shows that both the initialization and optimization of $\tau$ is crucial. A large initialization of $\tau$ requires many additional training steps to optimize to its correct value, whereas a low initialization results in training instability (Figure 5). For our negative mining model, we find $\tau=0.07$ to be a temperature initialization that is both stable and performant.
@@
Temperature in pretraining. We find that the mined negatives are essential for training stability. Figure 5 shows our pretraining run with a batch size of 128 images queries and 1024 text candidates. If randomized negatives are used instead of our chosen mined negatives, the optimizer tends to push $\tau$ very close to 00. This results in numerical instability, loss spikes, and eventually catastrophic failure. When our mined negatives are used in our pretraining, the temperature stabilizes at a very reasonable value of around $0.03$.
4.4 CtrlBench
To measure our model’s instruction following capabilities, we construct CtrlBench, an instruction-controlled retrieval benchmark. CtrlBench has a similar format to the Flicker30K (Plummer et al., 2016) test split. However, instead of each image having multiple valid captions, we instead provide an instruction from which the model can infer the most relevant text candidate. As each image has $5$ associated captions, a model that cannot utilize instructions can (in expectation) achieve at most $20\%$ R@1 on the benchmark. Therefore, CtrlBench tests both retrieval and instruction following capabilities. To construct CtrlBench, we sample a $1000$ test images from VisualGenome (Krishna et al., 2016). To create $5000$ instructions and text candidate pairs, we generate $5$ instructions for each image. We follow the same process for generating instructions as in Section 3.2. We remove duplicated captions from the dataset to prevent ambiguity in our text candidates. We filter CtrlBench to ensure that no images, instructions or retrieval candidates used during training are contained in the benchmark. We provide a discussion on CtrlBench’s motivation and design choices in Section 5.3 and samples in Appendix E.
Table 3: Performance on CtrlBench.
Model R@1 R@5 R@10
UniIR (Wei et al., 2023) 0.0 0.0 0.1
MagicLens (Zhang et al., 2024) 9.7 23.9 32.94
VLM2Vec (Jiang et al., 2024b) 24.0 49.7 61.1
ABC (ours) 39.7 69.1 78.9
Table 3 shows the performance of multimodal embedding models on CtrlBench. We compare with $2$ instruction fine-tuned CLIP derivatives: UniIR (Wei et al., 2023) and MagicLens (Zhang et al., 2024) as well as VLM2Vec (Jiang et al., 2024b), a concurrent work on adapting VLMs into multimodal embedding models. We find that neither of the CLIP-based architectures have above $20\%$ R@1 on CtrlBench, the performance that indicates that the model is non-trivially utilizing instructions. Conversely, the VLM architectures are better at instruction-controlled retrieval. Both ABC and VLM2Vec have R@1 above $20\%$.
4.5 VLM Backbone

### v2

@@ 4 Experiments
Table 1: Comparison of retrieval performance on MSCOCO Lin et al. (2015) and Flickr30K Plummer et al. (2016) datasets (Karpathy split). Best performance is bold, second best is underlined. * indicates a closed-weight model.
We evaluate two aspects of the model. As controlling a poor-quality embedding is useless, we first assess the quality of embeddings output by ABC. We evaluate our model on a variety of zero-shot retrieval and classification tasks (Section˜4.3). We study how variable resolution capabilities (Table˜3), the settings for $\tau$ when contrastively training a VLLM (Section˜4.5) and VLLM backbone choice (Section˜4.6) effect embedding quality. Secondly, we measure ABC’s ability to use natural language to control visual embeddings on several visual question answering tasks (Section˜4.7). Motivated by shortcomings in VQA benchmarks, we construct CtrlBench. A benchmark designed to require the model to jointly use the image and instruction to retrieve the most appropriate caption.
4.1 Training Settings
We initialize ABC using Qwen2-VL-7B  (Wang et al., 2024b). We train our negative mining model using only in-batch negatives with a batch size of $256$ for $1000$ steps. We set $\epsilon=0.95$ and sample 7 mined negatives for each image. Mined negatives are sampled across the pretraining dataset. We pretrain using batches of $512$ image queries and $4096$ text candidates sharded across 8 NVIDIA A100-SXM4-80GB GPUs (Qu et al., 2021) for 4000 steps. We use a LoRA adapter with a rank of $64$ and a fixed alpha of $128$. We limit the number of tokens output by the vision encoder to $512$ during training. For our optimizer, we use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $4\times 10^{-5}$, betas of $0.9$ and $0.999$ and a weight decay of $10^{-3}$. We warm-up for 3% of training steps and initialize the temperature $\tau$ as $7\times 10^{-2}$. In our instruction fine-tuning stage, we use a lower rank LoRA adapter. We set the rank and alpha to $16$ and $32$, respectively. Our instruction fine-tuning stage can be short, as our VLLM backbone is already instruction fine-tuned. Therefore, we only instruction fine-tune for $100$ steps. Each batch contains 128 unique images, with each image appearing four times, paired with a different instruction and a corresponding positive text candidate.
Modifications to save VRAM. Due to our use of LoRA Hu et al. (2021), the VRAM requirements for gradients and optimizer state is relatively low. Consequentially, the model activations used by autograd for the backward pass account for most of VRAM used during training. To address this, we make aggressive use of activation checkpointing, recomputing the activations of each decoder block during the backward pass. Furthermore, we modify the VLLM backbone to skip the logits and cross-entropy loss computation. With the large vocabulary size of modern LLMs, the output layer has become the most memory-intensive layer (Wijmans et al., 2024). As we only use the hidden state for calculating embeddings, the logits tensor is unnecessary. We find that this simple change saves up to 11 GB of memory per device, allowing us to use to a significantly larger batch size.
4.2 Inference Settings
When evaluating ABC we adjust the vision encoder to output a number of tokens corresponding to the resolution of the image. We rescale images to the nearest resolution that fits our vision encoder, which has a $14\times 14$ pixel patch size (Wang et al., 2024b). For very high resolution images, we cap the number of visual tokens used to encode an image to $1024$, down-scaling the image accordingly. When evaluating models that supports variable resolution, we vary their token budgets similarly (Jiang et al., 2024; Zhang et al., 2025). However, when we attempt to increase token budgets in CLIP models using the same method, we find that it degrades performance (Appendix˜G). We find that these models perform best when evaluated using their training resolution. Therefore, we do not scale the token budget when evaluating CLIP models. This includes instruction-finetuned CLIP models such as Uniir and MagicLens (Wei et al., 2023; Zhang et al., 2024).
@@ 4.2 Inference Settings
ImageNet-1K Russakovsky et al. (2015) 55.8 63.5 45.4 58.3 48.0 10.3 49.1 71.2
HatefulMemes Kiela et al. (2021) 51.1 51.7 47.2 56.4 49.0 49.6 51.0 52.1
VOC2007 Everingham et al. 50.7 52.4 64.3 66.2 51.6 52.1 74.6 81.4
SUN397 Xiao et al. (2010) 43.4 68.8 39.6 63.2 57.0 34.5 60.1 71.8
Place365 López-Cifuentes et al. (2020) 28.5 37.8 20.0 36.5 31.5 21.5 35.3 40.7
ImageNet-A Djolonga et al. (2020) 25.5 14.2 42.6 9.8 8.0 3.2 31.6 49.4
ImageNet-R Hendrycks et al. (2021) 75.6 83.0 75.0 66.2 70.9 39.7 66.2 86.8
ObjectNet Barbu et al. (2019) 43.4 51.4 40.3 32.2 31.6 20.6 49.2 67.7
Country-211 Radford et al. (2021) 19.2 16.8 14.2 11.3 6.2 2.5 9.3 18.5
@@
Table 2: Zero-shot classification results on MMEB (Jiang et al., 2024). We compare with pretrained CLIP models (Radford et al., 2021; Cherti et al., 2023a), instruction finetuned models derived from CLIP (Wei et al., 2023; Zhang et al., 2024) and other LLM backbone approaches (Li et al., 2023; Zhou et al., 2024).
Figure 4: (Left) Pixel distributions of benchmarks. ObjectNet (Barbu et al., 2019) contains significantly higer resolution than ImageNet (Russakovsky et al., 2015), which is downscaled. (Right) Accuracy on benchmarks when increasing the pixel count (tokens) used by vision encoder when creating embeddings.
4.3 Zero-shot Retrieval and Classification
@@
We first evaluate the quality of ABC embeddings, as controlling low-quality visual embeddings is useless. To test embedding quality, we use two zero-shot tasks: retrieval from a pool of candidates and image classification. For retrieval, we use MSCOCO (Lin et al., 2015) and Flickr30K (Plummer et al., 2016). For image classification, we use MMEB Jiang et al. (2024), which provides splits of many standard image classification tasks, such as ImageNet-1K (Russakovsky et al., 2015), VOC-2007 (Everingham et al., ), and ObjectNet (Barbu et al., 2019). We note that the classification labels in MMEB are short, often only one or two words. This is problematic for image-captioning models that have been trained on full sentences (Yu et al., 2022). To alleviate this issue, we use Radford et al. (2021)’s technique of embedding classification labels in a sentence template. We use “A photo of a {label}.” as our template for all classification evaluations.
In Table˜1, our model demonstrates strong image-to-text retrieval capabilities, achieving competitive performance to models that have been contrastively trained with hundreds of GPUs and massive batch sizes (Sun et al., 2023; Cherti et al., 2023a). We achieve the best performance in MSCOCO image-to-text retrieval. Comparatively, our text-to-image retrieval is weaker, as we under-perform both EVA-02-CLIP-E+ and OpenCLIP-G. Averaged across all classification tasks (Table˜2), we achieve $11.2\%$ better accuracy than the next-best model. On ImageNet-1K (Russakovsky et al., 2015), ABC has $7.7\%$ better zero-shot accuracy than OpenCLIP. ABC’s performance on established benchmarks such as MSCOCO and ImageNet-1K indicate that the pretrained model’s embeddings are useful. In Section˜4.7, we evaluate the fine-tuned adapter’s ability to embed images with text instructions into the pretrained model’s embedding space.
4.4 Image Resolution and Embedding Quality
# Negatives Image $\rightarrow$ Text Text $\rightarrow$ Image
0 50.4 35.7
3 58.4 44.0
6 59.9 44.8
Table 3: Top-1 retrieval performance on MSCOCO with varying number of mined hard negatives.
@@ 4.4 Image Resolution and Embedding Quality
Our VLLM backbone, Qwen2-VL-7B (Wang et al., 2024b), supports image inputs with variable resolution. This allows the user to effectively trade-off image resolution with inference speed by adjusting the number of tokens output by the vision encoder. We examine how this trade-off influences embedding quality for image classification. We find that performance on certain tasks, like ObjectNet (Barbu et al., 2019) and ImageNet-A (Djolonga et al., 2020), strongly correlate with the resolution used in the VLLM vision encoder (Figure˜4). On average, ObjectNet images have $13$ times more pixels than those from ImageNet-1K. When the number of tokens produced by the vision encoder is scaled up, we see a large improvement on ObjectNet, with accuracy increasing by 23.4%. However, lower resolution benchmarks like ImageNet-1K do not benefit nearly as much from scaling resolution. Notably, a smartphone camera takes photos with significantly higher resolution than images in ImageNet-1K or ObjectNet, by default (Apple Inc., 2024). This motivates the need for benchmarks containing larger resolution images. Existing low-resolution benchmarks may underestimate the capabilities of models that can natively utilize high-resolution images. Furthermore, we find that correctly classifying the “natural adversarial examples” of ImageNet-A (Djolonga et al., 2020) is largely a function of resolution. Our accuracy on ImageNet-A increases from only $12\%$ to $47.5\%$ simply by scaling the resolution used by the vision encoder during evaluation.
4.5 Mined Negatives and Temperature
@@
The setting of the temperature hyperparameter ($\tau$) is known to be crucial for contrastive training (Jia et al., 2021). However, $\tau$ is often treated as a fixed hyperparameter (Jiang et al., 2024; Wei et al., 2023). A large fixed setting of $\tau$ results in a significantly worse model (Wang & Liu, 2021), whereas a too small setting can result in training instability (Figure˜5). Therefore, we find that optimizing $\tau$ throughout the pretraining process is crucial. In shorter runs like training our negative mining model, we find that shows that both the initialization and optimization of $\tau$ is crucial. For our negative mining model, we find $\tau=0.07$ to be a temperature initialization that is both stable and performant.
@@
Mined negatives and temperature. In Table˜3, we ablate over the number of mined hard negatives used during pretraining. Each pretraining run consists of 500 steps with a batch size of 256. Introducing 6 mined hard negatives improves top-1 image to text retrieval by $9.5\%$ and top-1 text to image retrieval by $9.1\%$. Furthermore, we find that the mined negatives are essential for training stability. Figure˜5 shows our pretraining run with a batch size of 128 images queries and 1024 text candidates. We compare two runs: one using randomized negatives sampled uniformly randomly from our dataset and the other our mined hard negatives. We find that if randomized negatives are used instead of our mined negatives, the optimizer tends to push $\tau$ very close to 0. This results in numerical instability, loss spikes, and eventually catastrophic failure. When our mined negatives are used in our pretraining, the temperature stabilizes at a very reasonable value of around $0.03$.
4.6 VLLM Backbone